OpenStudy (anonymous):

prove the identity sin^2 acos^2=1/8(1-cos4a)

terenzreignz (terenzreignz):

Uhh... this one? \[\Large \sin^2(a) \cos^2(a) = \frac1{8}[1-\cos^4(a)]\]

OpenStudy (anonymous):

@terenzreignz Thank you for asking that. That's what I was wondering too :D @eyna_nuzrey

OpenStudy (anonymous):

@terenzreignz @genius12 nope it is cos4(a)

terenzreignz (terenzreignz):

Whoops\[\Large \sin^2(a) \cos^2(a) = \frac1{8}[1-\cos(4a)]\]

OpenStudy (anonymous):

@terenzreignz yes that's it!

terenzreignz (terenzreignz):

Well, there's not much we can do about the right-side, however, we can do something to the left... I suggest expanding cos(4a) using the double angle identities and the fact that 4a = 2(2a)

terenzreignz (terenzreignz):

whoops, got it backwards... I meant there's not much we can do about the *left* side, we can do something with the *right* sorry, my bad...

OpenStudy (anonymous):

ouh ok then?

terenzreignz (terenzreignz):

Yes, so... do that, expand cos(4a) and tell me what you get.

terenzreignz (terenzreignz):

To refresh you on the double angle identity for cosine, we have \[\large \cos(2\theta) = \cos^2(\theta) - \sin^2(\theta)\] With that in mind, what is \[\Large \cos(4a) = \cos[2(2a)]=\color{red}?\]

OpenStudy (anonymous):

erm just substitute the formula?

terenzreignz (terenzreignz):

Yes, substitute (2a) for \(\theta\) and you get...?

terenzreignz (terenzreignz):

By the way, our main objective is to somehow turn the right-side of the equation into the left side.

OpenStudy (anonymous):

\[\cos[2(\cos ^{2}a-\sin ^{2}a]\]

terenzreignz (terenzreignz):

No. First, expand this \[\Large \sin^2(a) \cos^2(a) = \frac1{8}[1-\color{red}{\cos(4a)}]\] Using this idea \[\large \cos(2\theta) = \cos^2(\theta) - \sin^2(\theta)\] With \[\Large \theta = 2a\]

OpenStudy (anonymous):

then?

terenzreignz (terenzreignz):

Then? You did incorrectly :) I said expand \[\Large \cos(4a)= \cos[2(2a)]=\color{red}?\] Using this principle \(\large \cos(2\theta) = \cos^2(\theta) - \sin^2(\theta)\)

OpenStudy (anonymous):

i don't understand which one i should expand?

OpenStudy (anonymous):

ouh i know!

terenzreignz (terenzreignz):

Cosine of 'twice something' regardless of what that something is, say, represent it with this triangle... \[\Large \cos(2\Delta)\] is just \[\Large \cos^2(\Delta) -\sin^2(\Delta)\] It just so happened that in this case, our 'triangle' is 2a. \[\Large \cos(4a) = \cos[2(2a)]\]

OpenStudy (anonymous):

ok i'm starting become blurr..

terenzreignz (terenzreignz):

What is cos(2a) ?

terenzreignz (terenzreignz):

Expand it, using the double-angle identity...

OpenStudy (bradely):

Step by step solution http://www.mathskey.com/question2answer/4229/prove-the-identity